OpenStudy (anonymous):

Find g(f(x)) if f(x) = x^2 + x + 3 and g(x) = 2x - 1. Not sure if I got it wrong or right but can you show me your solutions?

OpenStudy (anonymous):

g(f(x)) = g(x^2 + x + 3) = -1 + (x^2 + 2x + 3) = ?? Is my starting right or wrong?

OpenStudy (zzr0ck3r):

g(f(x)) = g(x^2 + x + 3) = -1 + 2(x^2 + x + 3)

OpenStudy (zzr0ck3r):

you did not distribute the 2 through for every ter,

OpenStudy (zzr0ck3r):

term*

OpenStudy (anonymous):

oh wait so you mean like -1 + (2x^2 + 2x + 6) ??

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

any place we see an x in 2x-1 we put x^2 + x +3 so 2(x^2+x+3)-1 2x^2+2x+6-1 2x^2+2x+5

OpenStudy (anonymous):

OH OK! thanks lots!

OpenStudy (zzr0ck3r):

np